### v2

# Mapping the Cosmic Gamma-ray Horizon: The 1CGH Catalogue of Fermi-LAT detections above 10 GeV

Abstract
[…]
keywords:
gamma-rays: galaxies, catalogues – radiation mechanisms: non-thermal
pubyear: 20252024pagerange: Mapping the Cosmic Gamma-ray Horizon: The 1CGH Catalogue of Fermi-LAT detections above 10 GeVB
1 Introduction
Blazars are a rare class of active galactic nuclei (AGN) characterised by relativistic jets pointing towards us, producing non-thermal emission that spans from radio to TeV $\gamma$-rays (Urry & Padovani, 1995; Blandford et al., 2019). Ranking among the most luminous and variable sources, blazars provide an unparalleled view into high-energy astrophysical processes, serving as cosmic beacons for studying the transparency of the universe to $\gamma$-rays.
The Fermi Large Area Telescope (Fermi-LAT, Atwood et al., 2009) has revolutionized our understanding of blazars, enabling us to collect data from tens of MeV to hundreds of GeV. Its vast energy coverage allows us to study the interaction of $\gamma$-rays with the Extragalactic Background Light (EBL), which leads to attenuation of the very high-energy (VHE) spectrum from distant sources (e.g. Ackermann et al., 2013; Fermi-LAT Collaboration et al., 2018). This process is driven by photon-photon annihilation via electron-positron pair production, $\rm\gamma_{VHE}+\gamma_{EBL}\rightarrow e^{+}+e^{-}$ (e.g. Nikishov, 1961; Gould & Schréder, 1967), turning $\gamma$-rays blazars into unique probes of the EBL density and evolution over cosmic time (e.g. Domínguez & Prada, 2013).
Direct EBL measurements are hindered by the presence of foregrounds, including zodiacal light (e.g. Finke et al., 2022). However, distant extragalactic $\gamma$-ray sources, like blazars, provide indirect probes to map the EBL density and distribution over cosmic time (e.g. Stecker et al., 1992). The EBL density in the far ultraviolet (FUV) to near-infrared (NIR) is largely driven by cumulative stellar activity along cosmic time (e.g. Hauser & Dwek, 2001; Stecker et al., 2016). Thus, measuring the EBL at a given redshift provides an alternative diagnostic to recover the star formation rate (SFR) from earlier epochs, reaching as far back as the Epoch of Reionization (e.g. Robertson et al., 2010) at $\sim$6.
Given the sensitivity of Fermi-LAT, observations are particularly well suited to measuring the EBL in the UV to optical bands, whereas the near-infrared (NIR) can still be constrained at lower redshift. For example, Fermi-LAT Collaboration et al. (2018) demonstrated how distant $\gamma$-ray sources can trace the EBL density to $\sim$3.0, providing unique insight into SFR close to the Epoch of Reionization (EoR).
The attenuation of $\gamma$-rays can be quantified by the EBL optical depth $\tau(E,z)$, which describes the cumulative probability of pair production and provides the relationship between observed and intrinsic flux:
$F_{observed}=F_{intrinsic}\times\exp^{-\tau_{(E,z)}}$ (1)
where a value of $\rm\tau_{(E,z)}=1$ implies that approximately 63% of the $\gamma$-ray flux at energy $E$from a source at redshift $z$- is absorbed. Measuring $\rm\tau_{(E,z)}$ is essential to understand the transparency of the universe to gamma-rays, and provides a framework for testing and validation of EBL models.
The attenuation of $\gamma$-rays by the EBL defines a statistical boundary for the farthest distance at which VHE photons of a given energy are likely to be detected, known as the Cosmic Gamma-ray Horizon (CGH). The concept of the CGH was initially explored by Nikishov (1961); Gould & Schréder (1967), while Fazio & Stecker (1970) -with the Fazio Stecker relation- advanced the discussion of $\gamma$-ray attenuation in context of a cosmological origin for the isotropic $\gamma$-ray background.
Interestingly, the VHE spectra of distant blazars can exhibit an unexpected hardening at the highest energies (e.g. Essey & Kusenko, 2010; Furniss et al., 2015). This phenomenon has led to alternative hypotheses regarding the propagation of VHE photons, including photon oscillations into axion-like particles (ALPs) (e.g. Galanti et al., 2019; Buehler et al., 2020; Abe et al., 2024) and potential Lorentz invariance violation (LIV) (e.g. Galanti et al., 2020; Abdalla et al., 2024). These mechanisms offer opportunities to probe fundamental physics and represent open questions in high-energy astrophysics. Therefore, a well-characterised sample of VHE emitters across a broad redshift range is essential, not only to probe EBL and SFR evolution more effectively, but also to investigate alternative photon propagation scenarios.
In this work, we use Fermi-LAT observations spanning the first 16 years of the mission (August 2008 to August 2024) to search for $\gamma$-ray signatures associated with a sample of blazars and blazar candidates. By focusing on sources detectable above 10 GeV, we aim to provide a catalogue where the EBL absorption is most pronounce and could be effectively measured in the highest energy bins.
To create a robust $\gamma$-ray sample suitable for studying EBL attenuation, we calculate the mean energy of the four highest-energy photons detected for each source. This approach offers a more stable and representative measure of the highest energy bin ($\rm E^{bin}_{max}$) detectable by Fermi-LAT, particularly when ranking sources that experience moderate to severe absorption, compared to using only the highest-energy photon ($\rm E_{max}$).
The First Cosmic Gamma-ray Horizon (1CGH) catalogue represents a significant step forward in describing the $\gamma$-ray horizon, and will contribute to our measurements of the EBL content and star formation rate across cosmic history.
[…]
In addition, we propose a selection criterion that focuses on sources above the $\rm\tau_{(E,z)}>0.1$ threshold, and build a subsample specifically targeted to the regime where at least 10% of the $\gamma$-ray flux is expected to be absorbed in the highest energy bin.
An essential component of this work is the extensive redshift review of the detected sources. Initially, nearly half of them lacked redshift estimates from the reference catalogues (5BZcat, 3HSP, 3FHL, 4LAC-DR3). To address this gap, we conducted a comprehensive literature search, covering nearly 70 publications. We gathered information on spectroscopic redshift, lower limit estimates, and uncertain redshift (including photometric estimates reported in 4LAC-DR3); and introduce a redshift quality flag to differentiate between those cases. This effort significantly improved the reliability of our redshift data, providing a stronger foundation for subsequent studies.
By providing a carefully curated sample of blazars with measurable EBL absorption, the 1CGH catalogue complements ongoing studies of TeV-detected sources (e.g. Gréaux et al., 2023) and aims to deepen the scientific discussion on cosmic $\gamma$-ray propagation.
2 Data Analysis and Methods
In this section, we discuss the selection of $\gamma$-ray candidates, the data analysis setup using the Fermi Science Tools, and the methods employed for cleaning and associating high-energy photons with sources detected above 10 GeV. We also outline the overall strategy used to ensure robust detections and photon associations, as reported in the 1CGH catalogue.
2.1 Selection of Gamma-ray Candidate Sources
To create a sample of sky positions and define our $\gamma$-ray targets, we combined several major blazar and AGN catalogues. We began by merging the ‘5th edition of the Roma-BZCAT Multi-Frequency Blazar Catalog’ (5BZcat, Massaro et al., 2015a) with the ‘3rd edition of the High Synchrotron Peak sample’ (3HSP, Chang et al., 2019). These catalogues were chosen for their extensive blazar coverage and reliability.
Both 5BZcat and 3HSP are advanced versions of multifrequency catalogues containing thousands of blazars and blazar candidates. These catalogues have been iteratively refined over time (e.g., refer to BZcat, 1WHSP, and 2WHSP catalogues for earlier versions Massaro et al., 2009; Arsioli et al., 2015; Chang et al., 2017), and represent significant repositories for the very high-energy community.
We also included the latest version (V3.4) of TeVcat (Wakely & Horan, 2008), which contains a follow-up list of VHE-detected sources; http://tevcat.uchicago.edu. Additionally, to improve the list of seed positions, we incorporate newly confirmed blazars and blazar candidates from recent literature, including Paliya et al. (2020); Maselli et al. (2013); Ighina et al. (2023, 2024).
We further include all sources classified as AGNs, blazars, and blazar candidates in the ‘Fourth Catalog of Active Galactic Nuclei Detected by the Fermi Large Area Telescope: Data Release 3’ (4LAC-DR3, Ajello et al., 2022), since many of those candidates are not covered by the 5BZcat and 3HSP catalogues.
To remove duplicate entries, we use TopCat (Taylor, 2005) for internal matching; For the 5BZcat and 3HSP samples, the R.A. and Dec. positions, corresponding to the optical and infrared counterparts respectively, were used. For TeVcat and 4LAC-DR3 sources, we used the associated counterpart position whenever available (e.g., radio, infrared, or optical), as these typically offer better astrometric precision compared to the $\gamma$-ray position. With this sample of unique sources, we searched for $\gamma$-ray associations in the 4FGL-DR4 catalogue (Ballet et al., 2023) and the 3FHL catalogue (Ajello et al., 2017), considering the 95% containment region reported for each source.
At this point, we have successfully listed unique seed positions recovered from the 5BZcat, 3HSP, TeVcat and 4LAC-DR3 catalogues; But still, many 4FGL-DR4 high-latitude sources ($|b|>10^{\circ}$) were not included in our initial sample of 10 GeV candidates. This is partly due to new $\gamma$-ray detections from DR4 that had not yet been incorporated into the latest 4LAC release. However, the main reason is that many of these missing sources are actually unassociated in the 4FGL-DR4 –lacking classification.
Therefore, in this work, we consider all unassociated high-latitude 4FGL-DR4 sources as valid 10 GeV candidates. Tracking detections above 10 GeV is essential, as it may motivate observational campaigns for their identification.
In the final sample of seed positions, we removed all sources flagged as extended and as $\gamma$-ray bursts (GRBs). Finally, we used the CLASS1 flag from the 4FGL-DR4 catalogue to exclude cases associated with galactic sources (e.g., pwn, lmb, hmb, msp, glc, bin, nov, spp, snr, Ballet et al., 2023). The resulting sample contains nearly 8200 seed positions, with 5023 flagged as blazars or blazar candidates. The remainder are unknown types of sources from 4FGL-DR4 and TeVcat. Given how the positional seeds were selected, the 1CGH catalogue includes unassociated $\gamma$-ray sources that require further observational efforts to determine their nature, i.e, to confirm whether they are extragalactic in origin.
Table 1: Power-law model parameters for the new $\gamma$-ray sources, with TS above 16, discovered in the 1CGH catalogue. A complete table with all 2791 1CGH sources is available in the online version of this paper (Note: A preliminary version is available in the Authors’ GitHub repository). The first three columns show the original source names, right ascension (R.A.), and declination (Dec.) in degrees (J2000), respectively. The fourth column lists redshifts from literature, marked with a ‘(?)’ flag for uncertain or photometric values, and with a ‘>’ symbol for lower limits. The power-law parameters regarding the 1CGH analysis are reported in the following columns, corresponding to the fit (Equation 2): the normalization ‘$\rm N_{0}$’ in units of ph/cm2/s/MeV, the photon spectral index ‘$\rm\Gamma$’, and the flux integrated over 10-800 GeV, given in units of ph/cm2/s. The pivot energy, ‘$\rm E_{0}$’, is fixed at 15 GeV for all sources. The ‘TS’ column indicates the Test Statistic value. The final column, $\rm E_{max}^{[GeV]}$, provides the highest energy photon associated with each 1CGH source, based on PASS8 UltraClean event-class, with PSF=0 events removed, i.e. evclass=512 and evtype=56 (see Section 2.4).
[…]
3HSPJ032356.5--010833 50.9855 --1.14272 0.391 --45.15 37.30$\pm$6.90 2.43$\pm$0.25 69.7$\pm$11.9 167 50.5
3HSPJ120711.5--174605 181.79787 --17.76830 >0.7 43.83 15.59$\pm$4.69 2.90$\pm$0.52 26.54$\pm$7.82 44.7 37.2
3HSPJ220451.0--181536 331.21379 --18.26011 0.26(?) --50.75 7.37$\pm$3.21 1.91$\pm$0.36 17.19$\pm$6.28 30.6 51.4
3HSPJ030103.7+344101 45.26558 34.68366 0.246 --21.00 8.95$\pm$3.37 2.29$\pm$0.51 17.46$\pm$5.96 30.5 32.1
3HSPJ132635.9+254958 201.64970 25.83300 0.698 82.02 4.46$\pm$2.33 1.62$\pm$0.35 12.88$\pm$4.97 28.9 116.8
3HSPJ090802.2--095937 137.00921 --9.99369 0.054 24.38 5.41$\pm$1.04 1.80$\pm$0.14 13.50$\pm$2.70 26.7 70.6
3HSPJ231023.4+311949 347.59737 31.33033 0.48(?) --26.77 5.30$\pm$0.31 1.85$\pm$0.11 12.88$\pm$1.15 26.1 51.8
3HSPJ155424.1+201125 238.60054 20.19038 0.273 47.76 10.22$\pm$4.11 2.09$\pm$0.47 21.69$\pm$8.78 25.9 138.6
3HSPJ101616.8+410812 154.07008 41.13680 0.27 55.32 4.04$\pm$2.22 1.77$\pm$0.39 10.39$\pm$4.41 24.1 164.2
3HSPJ231041.8--434734 347.67400 --43.79280 0.089 --63.75 3.86$\pm$0.30 1.71$\pm$0.10 10.37$\pm$0.94 23.6 48.7
3HSPJ094606.1+215138 146.52554 21.86066 0.489 47.73 7.27$\pm$2.83 2.94$\pm$0.77 12.34$\pm$4.89 22.0 20.3
3HSPJ091222.9--251825 138.09545 --25.30702 0.33(?) 15.62 7.76$\pm$3.04 2.45$\pm$0.50 14.41$\pm$6.00 21.8 52.1
3HSPJ141003.9+051557 212.51633 5.26605 0.544 61.21 6.99$\pm$3.33 2.26$\pm$0.51 13.79$\pm$5.65 21.6 68.0
3HSPJ061104.1+682956 92.76716 68.49905 0.5(?) 21.53 5.38$\pm$0.76 2.29$\pm$0.16 10.51$\pm$1.57 21.5 58.2
3HSPJ024115.5--304140 40.31454 --30.69447 0.3(?) --65.76 6.73$\pm$1.06 2.25$\pm$0.23 13.32$\pm$2.33 21.3 59.9
3HSPJ102523.0+040229 156.34595 4.04150 0.208 48.21 6.34$\pm$0.38 2.51$\pm$0.15 11.57$\pm$1.04 20.5 28.7
3HSPJ121038.3--252713 182.65970 --25.45383 0.47(?) 36.50 8.60$\pm$1.95 3.22$\pm$0.47 14.31$\pm$3.25 19.7 26.3
3HSPJ213004.8--563222 322.51987 --56.53947 -- --43.91 5.26$\pm$2.47 1.94$\pm$0.39 12.06$\pm$4.80 18.4 57.2
5BZQJ1353+0151 208.46492 1.86497 1.608 60.64 2.27$\pm$0.13 1.41$\pm$0.08 8.166$\pm$0.73 18.3 55.0
3HSPJ094502.0--044833 146.25845 --4.80938 0.43(?) 34.82 6.95$\pm$2.61 2.78$\pm$0.65 12.04$\pm$4.49 17.8 50.1
3HSPJ120543.3+582933 181.43029 58.49266 0.4(?) 57.63 5.15$\pm$2.22 3.44$\pm$0.98 8.539$\pm$3.68 17.8 24.5
3HSPJ213448.2--164205 323.70091 --16.70155 0.8(?) --43.51 7.97$\pm$3.48 3.63$\pm$1.19 13.23$\pm$5.74 17.3 29.5
3HSPJ054903.0--215001 87.26416 --21.83369 0.35(?) --23.05 5.12$\pm$2.45 2.46$\pm$0.66 9.495$\pm$4.17 17.2 66.5
5BZGJ1840--7709 280.16083 --77.15797 0.018 --25.80 5.97$\pm$2.57 2.44$\pm$0.52 11.10$\pm$4.54 16.8 29.8
3HSPJ064326.7+421418 100.86133 42.23852 0.089 16.47 7.31$\pm$3.28 2.81$\pm$0.72 12.62$\pm$5.34 16.5 68.0
3HSPJ100444.8+375212 151.18654 37.87000 0.44 53.58 4.52$\pm$0.65 3.86$\pm$0.47 7.580$\pm$1.11 16.3 13.7
3HSPJ151136.9--165326 227.90375 --16.89077 0.36(?) 34.38 8.68$\pm$3.59 3.02$\pm$0.82 14.63$\pm$5.93 16.3 30.8
3HSPJ104745.8+543741 161.94087 54.62813 0.54(?) 54.46 4.56$\pm$2.07 2.61$\pm$0.71 8.165$\pm$3.64 16.3 31.7
3HSPJ081941.8+053023 124.92433 5.50638 0.37(?) 22.11 7.02$\pm$3.26 2.80$\pm$0.75 12.14$\pm$5.52 16.2 17.0
3HSPJ062753.4--151957 96.97237 --15.33252 0.3102 --12.04 5.32$\pm$2.89 1.85$\pm$0.41 12.89$\pm$5.63 16.2 38.8
3HSPJ133326.0+623541 203.35820 62.59494 0.48(?) 53.86 1.69$\pm$0.10 1.43$\pm$0.08 5.954$\pm$0.35 16.0 93.2
2.2 Broadband Likelihood Analysis with Fermi-LAT Science Tools
To search for $\gamma$-ray signatures associated with the pre-selected seed positions, we performed a binned likelihood analysis in the 10-800 GeV band, covering the first 16 years of Fermi-LAT observations (August 2008 to August 2024). For the broadband analysis, we used the latest version of the Fermi Science Tools (V2.2.0) [^1] The Fermitools-Conda repository: https://github.com/fermi-lat/Fermitools-conda. The Fermi-LAT data analysis recommendations: https://fermi.gsfc.nasa.gov/ssc/data/analysis/., with Pass 8 event selection (P8R3, Atwood et al., 2013; Bruel et al., 2018), following the Fermi-LAT’s team recommendation for the identification of point-like sources.
We worked with FRONT+BACK source-class events (evtype=3 and evclass=128) and the instrument response function (IRF) P8R3-SOURCE-V3. For each analysis, we considered a region of interest (ROI) of 15 radius centred on the $\gamma$-ray seed positions. For modelling the background of point and extended sources, we adopted the 4FGL-DR4 v34 catalogue (gll-psc-v34.fit), along with isotropic and galactic-diffuse templates (iso-P8R3-SOURCE-V3-v1.txt and gll-iem-uw1216-v13.fits, respectively). The spectral parameters –normalization and photon index– were set free for all sources within 5 of the seed position.
A zenith angle cut of 105 was applied to avoid contamination from Earth’s limb $\gamma$-ray photons, which are produced by cosmic-ray interactions with the atmosphere. Using the gtmktime routine, we selected good time intervals when Fermi-LAT was operating in ‘science data-taking mode,’ by setting the flags DATA-QUAL$>$0 and LAT-CONFIG==1. Using the gtbin routine, we generated counts maps (CMAP) and counts cubes (CCUBE) with dimensions of $300\times 300$ and $200\times 200$ pixels at $0.1^{\circ}$/pixel, respectively. For the CCUBE, we used 20 equally-spaced logarithmic energy bins in the 10-800 GeV range. Each $\gamma$-ray candidate was modelled as a point-like source characterised by a power-law spectrum:
$\rm\hskip 85.0pt\frac{dN}{dE}\,\text{=}\,N_{0}\;\left(\frac{E}{E_{0}}\right)^{-\Gamma}\;,$ (2)
In this equation, $\rm N_{0}$ is the normalization constant (prefactor), given in units of [photons/cm2/s/MeV], representing the flux density at the pivot energy $\rm E_{0}$. The parameter $\rm\Gamma$ represents the photon spectral index, indicating the slope of the spectrum. To restrict the parameter space probed during the fitting process, we set a limit of 6.5 for the maximum $\rm\Gamma$ value. This limit is compatible with the largest spectral indices observed in the 3FHL catalogue, including all source types (i.e. Galactic, extragalactic, and unassociated/unknown sources, Ajello et al., 2017)
In our analysis, we adopt a low-energy threshold of 10 GeV –similar to the 3FHL catalogue– due to the improved angular resolution and lower background contamination compared to lower thresholds (Ajello et al., 2017). For the upper energy limit, we set a conservative threshold of 800 GeV, whereas 3FHL extended the analysis up to 2 TeV. This decision follows from the discussion in the 4FGL paper (Section 3.2 Abdollahi et al., 2020), which highlights that a broadband analysis reaching 1 TeV can introduce uncertainties in the energy flux estimate, particularly for hard-spectrum sources. Therefore, our choice for the high-energy threshold is meant to mitigate uncertainties regarding the spectral fitting of all sources in the ROI region.
2.3 Detection Threshold for Multifrequency-Selected Targets
The use of multifrequency seed positions to detect $\gamma$-ray sources has been successfully applied in numerous works (Arsioli & Chang, 2017; Arsioli et al., 2018; Arsioli & Polenta, 2018a; Arsioli et al., 2020) and has proven effective in uncovering extreme blazars close to the Fermi-LAT detection threshold. Using this approach, the analysis for each sky position is independent, investigating a single $\gamma$-ray candidate at a time. To handle all 8200 candidates efficiently, we relied on parallel High Performance Computing (HPC) resources provided by INCD-Portugal.
[…]
In addition, this relatively low selection threshold aims not only to highlight faint sources but also to capture those that undergo short-lived flare episodes, which might otherwise be missed over a long integration period due to signal dilution (Arsioli & Polenta, 2018b). As the true signal (i.e. the photon counts from a transient flare) remains constant while the background noise accumulates, a lower TS threshold allows us to include those valuable cases. In short time windows, the significance of these flaring events would be higher, enabling the study of the effects of the EBL absorption. Although we did not perform light-curve analyses in this work, the 1CGH catalogue can serve as a basis for future time-resolved studies.
To assess the robustness of our detection threshold, we estimated the spurious detection rate in our analysis setup. In a chi-squared ($\chi^{2}$) distribution with two degrees of freedom (d.o.f), the cumulative probability $P(TS>TS_{threshold})$ represents the likelihood of obtaining a test statistic above the threshold purely by chance. Using TS=12, the spurious detection probability is calculated as $\rm p\_value=1-\chi^{2}\times cdf(TS=12,\ d.o.f=2)\approx 0.0025$, where $cdf()$ is the cumulative distribution function. Given that 8200 seed positions were analysed, the expected number of spurious detections in the pre-selection phase is approximately $\rm 8200\times 0.0025=20.5$. This value represents an upper limit on contamination.
In next section, we discuss an additional selection criteria –beyond the selection based on TS level alone– to avoid spurious detections and improve reliability of the final catalogue.
2.4 Additional Selection Criteria: Refining the 1CGH Sample
As a result of the likelihood analysis, we pre-selected 3004 sources with TS > 12, which are used as a base for building the final sample. Those sources will still go through an additional selection step that involves looking at the $\gamma$-ray events associated with each source. In this stage, it is essential to consider the different event types in the dataset and also to identify and remove potential sources of contamination from the event sample.
In the Fermi-LAT PASS8 release (Atwood et al., 2013), each event is assigned a Point Spread Function (PSF) type, which indicates the quality of the reconstructed direction. Events are categorized into four quartiles: PSF0 (lowest quality, evtype=4), PSF1 (evtype=8), PSF2 (evtype=16), and PSF3 (highest quality, evtype=32). The PSF 68% containment radius varies by energy and event type. For instance, at 30 GeV, the containment radius for PSF1 events is approximately 0.12, as described in the PASS8 documentation [^4] Fermi-LAT PASS8 PSF: https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm.
Using the gtselect routine, we created a subsample of the all-sky Fermi-LAT data, selecting only events with energy greater than 10 GeV. We removed low-quality events by excluding PSF0 events, which represent the quartile with the poorest reconstruction accuracy. Next, we used the gtmktime routine to select events during good time intervals (see Section 2.2). After this step, we were left with a photon sample containing PSF1-2-3 source-type events (i.e., evclass=128 and evtype=56), with a maximum zenith angle of 105, all recorded under ‘science data-taking mode’.
Before recovering the $\gamma$-ray events associated with the pre-selected sources, we applied an additional layer of data cleaning to improve the quality of the photon sample. Specifically, we removed events associated with solar emissions and $\gamma$-ray bursts (GRBs). Indeed the solar disk has recently been confirmed as a source of very high-energy $\gamma$-rays (Linden et al., 2022; Albert et al., 2023; Arsioli & Orlando, 2024); Therefore, we removed all events that might be linked to solar emissions by assuming an association radius of 0.8555Considering the solar disk has an angular radius of $\sim$0.26, and the characteristic position uncertainty for PSF1 events at 10 GeV is around 0.18, a 0.8 radius cut safely encompasses and removes solar disk contamination..
To identify and remove events related to GRBs, we used the Fermi GBM Burst Catalogue  [^6] The continuously updated GBM Burst Catalogue, also referred to as ’The FERMIGBRST database’, is available at https://heasarc.gsfc.nasa.gov/W3Browse/fermi/fermigbrst.html. (von Kienlin et al., 2020). We removed events recorded within 0.8 of GRB locations, also considering a time window of 30 minutes before and 10 hours after each GRB.
The characteristic position uncertainty of PSF1 events at 30 GeV was used as a basis for defining a cross-matching radius of 0.12 between the pre-selected candidates and the clean photon sample. This conservative radius ensures robustness when associating higher-energy events, particularly given the larger PSF1 size at the lowest energy (approximately 0.18 at 10 GeV).
As an intermediary selection layer of the 1CGH catalogue, we retained only pre-selected sources (TS>12) that were associated with at least four high-energy events from the clean photon sample, matching within 0.12. This approach follows similar criteria used in building the 3FHL sample (Ajello et al., 2017), which required a minimum of four associated photons for source acceptance (as predicted by the adjusted model, i.e., $N_{pred}\geq 4$). This procedure removed 213 sources ($\rm\sim 7\%$) from the 3004 pre-selected ones. We are left with a final sample of 2791 sources, and –given the selection criteria described above– the expected level of spurious detections should be lower than the 0.68% figure (i.e., 20.5/3004) presented in Section 2.3.
3 The First Cosmic Gamma-ray Horizon Catalogue (1CGH)
We present the First Cosmic Gamma-ray Horizon catalogue (1CGH), which lists 2791 blazars and blazar candidates detected with a TS greater than 12 in the 10-800 GeV energy range, integrated over 16 years of Fermi-LAT observations. For detailed information on the catalogue’s metadata, refer to Table 2 in Appendix A.
The 1CGH catalogue includes 62 $\gamma$-ray detections never reported in earlier Fermi-LAT catalogue releases (1FGL, 2FGL, 3FGL, 4FGL, 2FHL and 3FHL), representing new $\gamma$-ray sources. These newly identified sources are mostly high synchrotron peak blazars from the 3HSP catalogue (Chang et al., 2019), and could be especially significant as candidates for VHE observations with the upcoming Cherenkov Telescope Array Observatory (CTAO, Cherenkov Telescope Array Consortium et al., 2019; Acharya et al., 2013). The 3HSP catalogue was designed to identify VHE targets for CTAO, and here we confirm its potential.
Table 1 presents a selection of newly detected sources with TS above 16, including the seed-source name, right ascension (R.A.), declination (Dec.), redshift, and power-law model parameters associated with each new $\gamma$-ray detection [^7] The full version of the First Cosmic Gamma-ray Horizon catalogue will be available in public data repositories such as Vizier https://vizier.cds.unistra.fr/viz-bin/VizieR and Author’s GitHub..
Figure 1: The distribution of detection significance ($\sigma$) for the 1CGH and 3FHL samples (respectively, blue and red-dashed lines).
In Fig. 1 we show the distribution of detection significance for the entire 1CHG and 3FHL catalogues. This representation is meant as a qualitative overview on both samples, keeping in mind that the 1CGH is focus on seed-positions of blazars and blazar-candidates, while the 3FHL includes detections of all source types. The number of detections in 1CGH shows consistent growth along the significance-bins (i.e. following a trend similar to 3FHL), and includes an extra bin covering the 3<$\sigma$<4 range.
We should note that, in the full 1CGH catalogue, 26 sources [^8] Comprising 3 unassociated sources, 4 BL Lacs, 8 Flat Spectrum Radio Quasars (FSRQs), and 11 blazars of unknown type. reached the spectral index limit of $\Gamma=6.5$ constrained by our likelihood analysis setup (see Section 2.2). Moreover, 27 1CGH sources have counterparts in 3FHL where spectral curvature has been identified [^9] To help identify such cases, we include the Curvature_3FHL column to our final catalogue, which tracks whether spectral curvature was identified in 3FHL, and might affect broadband flux estimates.. For these 53 (26+27) cases, the derived fluxes should be interpreted as upper limits. While this affects a small fraction of the sample, the vast majority of sources exhibit well-constrained spectral fits, reinforcing the robustness of the 1CGH catalogue for high-energy studies.
3.1 Extensive Redshift Review
In addition to listing detections above 10 GeV, we conducted an extensive literature review to improve the redshift characterization of our sample, which is crucial for investigating the cosmic $\gamma$-ray horizon. Based on the latest studies, we revised and updated redshift information, tracked the origin and quality of the data, and incorporated new redshifts from multiple optical campaigns dedicated to characterizing blazars, 3FHL and 4FGL sources. To track redshift quality, we added a column named ‘z-flag’ to the catalogue, with values: (1) for robust spectroscopic redshift; (2) for photometric estimates or uncertain values reported in literature; and (3) for lower-limit redshift values.
In the 1CGH catalogue, we added a column named ‘z-origin’ to track the origin of the redshift values reported in the ‘z’ column. Redshift information was gathered from several major blazar and $\gamma$-ray catalogues, including 5BZcat (Massaro et al., 2015a), 3HSPcat (Chang et al., 2019), 4LACdr3 (Ajello et al., 2022), and TeVcat (Wakely & Horan, 2008).
Additionally, redshift values were extracted from dedicated observational campaigns, such as Boyle90 (Boyle et al., 1990), Appenzeller98 (Appenzeller et al., 1998); Sbarufatti05 (Sbarufatti et al., 2005); Sbarufatti06 (Sbarufatti et al., 2006); Sbarufatti09 (Sbarufatti et al., 2009); Maisner10 (Meisner & Romani, 2010); Pulido10 (Acosta-Pulido et al., 2010); Shaw12 (Shaw et al., 2012); Furniss13 (Furniss et al., 2013); Landoni13 (Landoni et al., 2013); Maselli13 (Maselli et al., 2013); Rovero13 (Rovero et al., 2013); Sadrinelli13 (Sandrinelli et al., 2013); Shaw13 (Shaw et al., 2013); Paggi14 (Paggi et al., 2014); Ricci15 (Ricci et al., 2015); AlvarezCrespo16a-b-c (Álvarez Crespo et al., 2016a, b, c); Kaur16 (Kaur et al., 2017); Rovero16 (Rovero et al., 2015, 2016); Juanita17 (Torres-Zafra et al., 2018); Paiano17 (Paiano et al., 2017c, a, b); Gabanyi18 (Gabányi et al., 2018); Landoni18 (Landoni et al., 2015, 2018); Massaro18 (Massaro et al., 2014, 2015b, 2016); Mishra18 (Mishra et al., 2018); Balmaverde19 (Balmaverde et al., 2020); Caccianiga19 (Caccianiga et al., 2019); Johnson19 (Johnson et al., 2019); Desai19 (Desai et al., 2019); Menezes19 (de Menezes et al., 2019, 2020); Marchesini19 (Marchesini et al., 2019); Paiano19 (Paiano et al., 2019); Belladitta20 (Belladitta et al., 2020); Goldoni20 (Goldoni et al., 2015, 2021); Landoni20101010ZBLLac Database: “A spectroscopic Library of BL Lac objects” at https://web.oapd.inaf.it/zbllac/ (Landoni et al., 2020); Pena-Herazo20 (Peña-Herazo et al., 2020, 2019, 2017); Paiano20 (Paiano et al., 2019, 2020); Raiteri20 (Raiteri et al., 2020); B.Gonzalez21 (Becerra González et al., 2021); Paliya21 (Paliya et al., 2021); Pena-Herazo21 (Peña-Herazo et al., 2021b, a); Rajagopal21 (Rajagopal et al., 2021); Foschini22 (Foschini et al., 2022); Kasai22 (Kasai et al., 2023b, a); OlmoGarcia22 (Olmo-García et al., 2022); Rajagopal22 (Rajagopal et al., 2023); Paiano23 (Paiano et al., 2023); dAmmando24 (D’Ammando et al., 2012, 2024); Sarira24 (Sahu et al., 2024), DESI-EDR (DESI Collaboration et al., 2024), and AlvarezCrespo25 (Álvarez Crespo et al., 2025). The numerous dedicated optical observation campaigns highlight the ongoing high demand for the characterization of blazars and blazar candidates, particularly in connection with their $\gamma$-ray counterparts.
In Figure 2, we show the redshift distribution for the 1CGH subsample with z-flag=1 (i.e. robust spectroscopic redshift) and for the entire 3FHL sample. The comparison reveals that the number of sources detected above 10 GeV has improved along redshift relative to 3FHL. In addition, our redshift characterization aims to complement and refine previous efforts (e.g., 3FHL and 4LAC-DR3) by incorporating redshift quality flags and systematically tracking the reference from which each redshift value was obtained [^11] Please refer to Sec. 5.2 of Ajello et al. (2020) regarding redshift contamination in 4LAC, given the lack of flags to differentiate between spectroscopic and photometric redshift values..
Figure 2: The redshift distribution for the 1CGH and 3FHL samples. For the 1CGH, the histogram only include sources with spectroscopic redshift determination (z-flag=1); for the 3FHL catalogue, we include the entire sample.
Extensive redshift characterization is key for samples used to study the CGH, as it helps to mitigate possible biases caused by missing or uncertain redshift values. The best way to reduce uncertainties when measuring the EBL density evolution is to increase the number of sources with robust redshift determinations. In particular, identifying absorbed sources above redshift 2 would significantly improve EBL measurement at a regime that impacts SFR density estimates near and within the EoR, at $\sim$6-7 Fermi-LAT Collaboration et al. (2018). Among the most distant 1CGH sources (at $\rm 2.4\leq z\leq 2.9$) are: 5BZQJ0601-7036, 5BZQJ0228-5546, 5BZQJ1344-1723, 5BZQJ1748+3404, 5BZQJ1441-1523, 5BZQJ0912+4126, 5BZQJ1345+4452 and 5BZBJ0022+0608, which could contribute to studies of the EoR.
With improved precision in EBL density measurements, we can potentially disentangle the contributions of star formation and AGN to the EBL build-up at high redshift, turning blazars into an effective tool to understand the role played by AGNs during the EoR (e.g., D’Silva et al., 2023).
3.2 The Cosmic Gamma-ray Horizon Plot
Figure 3: The cosmic gamma-ray horizon, showing the highest-energy photon versus redshift, based on Fermi-LAT ultra-clean events with PSF0 events removed (evclass=512 and evtype=56). Sources detected above 10 GeV are represented, with robust (spectroscopic) redshifts marked as filled circles; lower-limit redshifts are represented by right arrows, with the reported redshift value corresponding to the center of the arrow
; and uncertain redshifts (photometric or reported as doubtful) indicated by blue crosses. The EBL optical depth $\rm\tau_{(E,z)}=1.0$ is depicted for different models: Finke et al. (2010) (blue dashed line), Domínguez et al. (2011) (black dotted line), Franceschini & Rodighiero (2017) (grey line), and Saldana-Lopez et al. (2021) (magenta dot-dashed line). For better visualization, the photon energy of sources with robust redshift is colour-coded. Note that approximately one-third of the 1CGH sources lack assigned redshift and are therefore not included in this plot.
Plotting the highest-energy photons from the 1CGH catalogue against redshift reveals that the universe is indeed opaque to $\gamma$-rays, as predicted by Nikishov (1961); Gould & Schréder (1967). As previously discussed, the energy vs. redshift relationship, where the opacity $\rm\tau(E,z)=1.0$, defines what is known as the ‘Cosmic Gamma-Ray Horizon’. Beyond this horizon, high-energy photons are severely attenuated by the EBL, rendering the universe effectively opaque to $\gamma$-rays (e.g. Domínguez et al., 2013).
Figure 3 shows the highest-energy photon versus redshift for the 1CGH sources, along with the predicted gamma-ray horizon according to the main EBL models (e.g. Finke et al., 2010; Domínguez et al., 2011; Franceschini & Rodighiero, 2017; Saldana-Lopez et al., 2021). Different markers are used to represent robust (spectroscopic), photometric/uncertain, and lower limit redshift values to emphasize the importance of accurate redshift determination.
Figure 3 provides an unprecedented view of the redshift characterization of sources used to study the CGH. It highlights the significance of reaching an extensive redshift description of the entire 1CGH sample and aims to further motivate the community’s ongoing efforts in the optical characterization of these rare sources (see Section 3.1).
Here we should draw attention to the number of 1CGH sources which lack a proper redshift characterization: Of the total 2791 1CGH sources, 1062 (38.2%) have spectroscopically measured redshift values, 855 (30.6%) have uncertain/photometric values, 210 (7.5%) have lower limit values, and 664 (23.7%) have no available redshift information. Therefore, 1729 ($\sim$72%) of the 1CGH catalogue would benefit from dedicated observation campaigns to improve the overall redshift description.
3.3 Identification of Sources with Detectable EBL Absorption
Figure 4: The cosmic gamma-ray horizon, showing the ‘largest energy bin detectable with Fermi-LAT’ versus redshift. The ‘largest energy bin’ is calculated as the mean energy of the four highest-energy source-type events associated with each source, excluding PSF0 events (i.e., evclass=128 and evtype=56). The opacity regimes for EBL optical depths $\rm\tau_{(E,z)}=1.0$ and $\rm\tau_{(E,z)}=0.1$ are represented by grey dashed and magenta dot-dashed lines, respectively, corresponding to the Saldana-Lopez model (Saldana-Lopez et al., 2021). This plot highlights sources in the $\rm\tau_{(E,z)}>0.1$ regime where absorption may be detectable at the highest energy bin observed with Fermi-LAT, while sources in the $\tau<0.1$ regime are represented by grey markers. Different markers indicate the quality of the redshift determination: robust (spectroscopic) redshifts are shown as filled circles; lower-limit redshifts are represented by right arrows, with the reported redshift value corresponding to the center of the arrow; and uncertain redshifts (photometric or unknown quality) are represented by blue crosses. Approximately one-third of the 1CGH sources lack redshift assignments and are therefore not included in this plot. Note: the $\rm\tau_{(E,z)}$ values regarding the largest energy bin are listed in the 1CGH catalogue to help identify interesting targets
To measure the EBL optical depth, one needs to build the $\gamma$-ray spectrum and compare the observed flux with the expected flux as a function of energy. For sources observed with Fermi-LAT, the $\gamma$-ray spectrum typically spans from tens of MeV to hundreds of GeV, and the flux can only be resolved for a few energy bins ($\rm E^{bin}$) (i.e. due to sensitivity limitations, Fermi-LAT usually cannot achieve high spectral resolution).
If we look for the most interesting sources for studying the cosmic gamma-ray horizon, we should focus on estimating the largest energy bin ($\rm E^{bin}_{max}$) where a source can still be detected. Any source with at least one energy bin that experiences a detectable degree of EBL attenuation is relevant to measure $\rm\tau_{(E,z)}$ values.
To emphasize sources that could be used to measure $\rm\tau_{(E,z)}$, we adopt a selection method based on the (z,$\rm E^{bin}_{max}$) position in the energy versus redshift plane. We calculate $\rm E^{bin}_{max}$ as the mean energy of the four highest-energy photons associated with a source, to create a robust estimate of the highest-energy bin. A similar requirement for the minimum photon counts was applied in the 3FHL catalogue (Ajello et al., 2017) to set a lower acceptance level for robust source detections. When calculating $\rm E^{bin}_{max}$, we use the clean photon sample (as described in Section 2.4), considering source-type events and excluding PSF0 events (i.e., evclass=128 and evtype=56).
In Figure 4, we illustrate the energy versus redshift relation and use the Saldana-Lopez EBL model Saldana-Lopez et al. (2021) as a reference framework to identify sources where gamma-ray absorption might be detectable. The Saldana-Lopez model offers an empirical determination of the evolving EBL spectrum and its uncertainties up to $\sim$6; it is based on galaxy counts and multifrequency data (from UV to far-IR) covering the Hubble Space Telescope’s CANDELS survey (Grogin et al., 2011), and was designed to minimise uncertainties especially at higher redshift.
Our ultimate goal, however, is to highlight a sample in which $\tau_{(E,z)}$ values can be derived from the data, independently of any specific EBL model. The estimates of Saldana-Lopez et al. (2021) are similar to those by Finke et al. (2010) and both are representative of the most severe attenuation models among the ones considered.
We use an EBL optical depth of $\rm\tau_{(E,z)}=0.1$ as a lower limit to introduce an absorption flag (ABS-flag), indicating whether spectral absorption might be detectable at the highest energy bins. For cases where $\rm E^{bin}_{max}$ lies in the $\rm\tau_{(E,z)}>0.1$ regime, the largest energy bin observed with Fermi-LAT likely experiences a detectable absorption greater than 10% (i.e., >($\rm 1-e^{-0.1}$)). These sources can be used to investigate the opacity along redshift, and have an ABS-flag set to ‘1’.
For sources with $\rm E^{bin}_{max}$ in the $\rm\tau_{(E,z)}<0.1$ regime, the Fermi-LAT spectrum is likely unaffected by EBL attenuation, and these sources have an ABS-flag set to ‘0’ in the 1CGH catalogue. These sources can be used to study the intrinsic $\gamma$-ray spectrum from blazars, helping to refine assumptions regarding their intrinsic flux.
The $\rm\tau_{(E,z)}$ values  –calculated using the Saldana-Lopez EBL model at $\rm E^{bin}_{max}$ Saldana-Lopez et al. (2021)– are listed under the column ‘$\rm\tau_{Ebin}$’. Additionally, the absorption fraction ($\rm 1-e^{-\tau_{(E,z)}}$) is recorded in the ‘$\rm Abs_{Ebin}$’ column, which can be used to create subsamples based on different absorption thresholds. Notably, for sources with multiple energy bins experiencing detectable absorption, a single source can provide $\rm\tau_{(E,z)}$ estimates for multiple energy levels at a given redshift.
3.4 Best Candidates for Optical Observation Campaigns
Following our extensive review of redshift information (sec. 3.1), Appendix B highlights the high galactic latitude sources (|b| > 10) that are currently the best candidates for optical observations. We have identified two main types of sources to prioritize, which will further improve the catalogue and the science described above:
  • 1CGH sources that currently lack redshift information (z-flag = 0) but already have an optical or radio association are prioritised based on $\rm E^{bin}_{max}$, the highest energy bin detected with Fermi-LAT. These sources are particularly significant due to the presence of a clear counterpart for optical follow-up and their potential proximity to the $\rm\tau\sim 1$ horizon. Table 31 highlights 42 of these cases.
  • 1CGH sources with lower limit redshifts or photometric/uncertain redshifts (z-flag = 2 or 3) are prioritised based on their calculated opacity ($\rm\tau_{Ebin}$), using $\rm E^{bin}_{max}$ values. These sources could be experiencing significant $\gamma$-ray absorption, potentially detectable across multiple energy bins, and thus could provide valuable $\rm\tau{{}_{(E)}}$ measurements across different energies at a given redshift. Table 42 highlight 43 of these cases.
Any candidate for follow-up could also be considered in light of the redshift upper limits proposed by Domínguez et al. (2024), which use EBL attenuation to constrain the distance of $\gamma$-ray blazars.
4 Summary and Conclusions
In this work, we presented the First Cosmic Gamma-ray Horizon (1CGH) catalogue, which lists all-sky Fermi-LAT detections above 10 GeV, compiled from 16 years of observations. The catalogue includes 2791 sources, with 62 representing new $\gamma$-ray detections, and to date, it represents an advanced sample to study the transparency of the universe to VHE photons.
We significantly improved the redshift characterization of our catalogue by incorporating new spectroscopic redshift estimates whenever possible. By meticulously reviewing nearly 70 dedicated observational campaigns, we reduced a substantial portion of the redshift knowledge-gap, categorising redshift reliability through our z-flag system.
The 1CGH catalogue highlights sources experiencing varying degrees of $\gamma$-ray attenuation, and will contribute to a better determination of the EBL’s density and its evolution over cosmic time. Using the Saldana-Lopez EBL model as a reference framework, we identified sources with $\gamma$-ray absorption above the $\rm\tau_{(E,z)}>0.1$ limit, creating a targeted sample where absorption is likely measurable.
In addition, we introduced a complementary data-exploration method by calculating the mean energy of the four highest-energy photons $\rm E^{bin}_{max}$, rather than focusing solely on the highest-energy event. This provides a more stable and statistically reliable measure of the characteristic ‘largest energy level’ observed for each source.
We have identified sources that should be prioritised for follow-up optical observation campaigns. These include: (i) sources that lack redshift information but have a clear optical or radio association, prioritised based on the highest $\rm E^{bin}_{max}$ values; and (ii) sources with lower limit or uncertain redshift estimates that could benefit from further spectroscopic exploration, prioritised by their current $\tau$ values at $\rm E^{bin}_{max}$. These targets will not only help refine the redshift description of the catalogue but also improve the accuracy of EBL measurements at higher redshift, thus reducing uncertainties in SFR density determination during the EoR, at $\sim$6-7.
Acknowledgements
The authors thank the anonymous Referee for the careful review and helpful comments that improved the quality -and impact- of our work. BA thanks the Institute of Astrophysics and Space Sciences (IA) at the University of Lisbon for their ongoing support. BA is currently a Marie Skłodowska-Curie Postdoctoral Fellow at IA, funded by the European Union’s Horizon 2020 research and innovation program under the MSCA agreement No. 101066981. BA also acknowledges the support from ‘Fundação para a Ciência e a Tecnologia’ (FCT) through the research grant UIDP/04434/2020 (DOI: 10.54499/UIDP/04434/2020) in support of general IA activities. This work was produced with the support of the "National Distributed Computing Infrastructure" (INCD), Portugal (https://www.incd.pt/), funded by FCT and FEDER under project 01/SAICT/2016 No. 022153. We acknowledge FCT’s support through the A1 Computation Call (project’s doi: 10.54499/2023.09549.CPCA.A1) for the allocation of HPC resources at INCD. We are grateful to the entire Fermi-LAT collaboration for maintaining a publicly accessible mission database, enabling discoveries across the entire $\gamma$-ray community. We also used archival data and bibliographic information from the NASA-IPAC Extragalactic Database (NED), and data/software facilities maintained by the Space Science Data Center (SSDC) from the Italian Space Agency. We extend our thanks to the TeVCat team, especially Deirdre Horan and Scott Wakely, for their continuous work on VHE detections, as well as NASA and NSF for their support of these follow-up efforts.
Data Availability
All catalogues used to build the sample of $\gamma$-ray candidates are publicly available and cited within this work. The 1CGH catalogue will be made publicly available through VizieR at https://vizier.cds.unistra.fr/, and GitHub. The Fermi-LAT database and the Science Tools used to build the 1CGH catalogue are also publicly accessible at https://fermi.gsfc.nasa.gov/ssc/data/access/.
Appendix A Column Descriptions of the 1CGH Catalogue
The 1CGH catalogue contains detailed metadata for 2791 sources detected above 10 GeV based on 16 years of Fermi-LAT observations. Table 2 provides a description of each column, detailing its content.
Table 2: Column descriptions of the 1CGH catalogue. The table lists the column names, units where applicable, and detailed explanations of their contents.
Column Name Description
1CGH_name Source identifier, formatted as 1CGHJ0123+0123, based on J2000 coordinates. Includes four digits each for R.A. and Dec. For associated sources, the coordinates are from the counterpart; for unassociated sources, we follow 4FGL-DR4 astrometry.
RA_1CGH Right Ascension (J2000) in degrees.
DEC_1CGH Declination (J2000) in degrees.
Counterpart_name Name of the source from catalogues such as 5BZcat, 3HSP, 4FGL-DR4, or others.
N0[1E-16] Normalization constant of the power-law fit in units of photons/cm2/s/MeV at the pivot energy.
N0_err[1E-16] Uncertainty in the normalization constant.
Index Photon spectral index of the power-law model.
Index_err Uncertainty in the photon spectral index.
Flux[1E-12] Integrated photon flux (10–800 GeV) in units of photons/cm2/s, estimated with a power-law fit.
Flux_err[1E-12] Uncertainty in the integrated photon flux.
TS Test Statistic value for the source detection.
sigma The significance for the source detection.
z Redshift of the source.
zflag Redshift reliability flag: (1) spectroscopic, (2) photometric/uncertain, (3) lower limit.
z_origin Reference for the redshift, as described in the text. For cases with no corresponding redshift value, it tracks sources where spectroscopic observation has been attempted.
Nph_128 Number of photons associated (Source_type events, evclass=128).
Emax_128[GeV] Maximum photon energy detected (evclass=128).
Nph_512 Number of photons associated (UltraClean_type events, evclass=512).
Emax_512[GeV] Maximum photon energy detected (evclass=512).
Ebin_max[GeV] Mean energy of the four highest-energy photons (evclass=128).
TransmissionEbin Flux transmission factor at $\rm E^{bin}_{max}$.
AbsEbin Flux absorption factor at $\rm E^{bin}_{max}$, calculated as ($1-$ Transmission).
tauEbin EBL optical depth at $\rm E^{bin}_{max}$.
Abs_flag Absorption flag: (1) $\tau>0.1$ (moderate to severe absorption), (0) $\tau\leq 0.1$.
GAL_LAT Galactic latitude in degrees.
BZcat Identifier in the 5BZcat catalogue, if available.
z_3HSP Redshift from the 3HSP catalogue, if available.
zflag_3HSP Redshift flag from the 3HSP catalogue.
4FGL_counter_name Counterpart name in 4FGL-DR4.
TEVCAT_FLAG_4FGLdr4 Flag for TeVCat association, as per 4FGL-DR4.
ASSOC_TEV_4FGLdr4 TeV association in 4FGL-DR4.
CLASS1_4FGLdr4 Classification in 4FGL-DR4.
ASSOC1_4FGLdr4 Counterpart name in 4FGL-DR4.
Name_4LACdr3 Counterpart name in 4LAC-DR3.
z_4LACdr3_1 Redshift as reported in 4LAC-DR3.
Name_3FHL Counterpart name in 3FHL.
z_3FHL Redshift as reported in 3FHL.
Curvature_3FHL Detectable spectral curvature in 3FHL, (1)yes, (2)no, (-) unknown. Sources marked as "unknown" are out of the 3FHL catalogue.
Appendix B Long Tables: Best Candidates for optical Observations
In this section, we present tables highlighting the best candidates for optical follow-up observations from the 1CGH catalogue. These sources have been selected based on their potential to provide insights about the CGH. Two categories of sources are prioritized: Table 31 those without redshift information but with clear optical, infrared, or radio associations, and Table 42 those with uncertain or lower-limit redshifts where significant $\gamma$-ray absorption is predicted. These candidates will refine the redshift completeness of the 1CGH catalogue. Each source selection is based -respectively- on its highest-energy bin ($\rm E^{bin}_{max}$), and on its calculated optical depth $\rm\tau_{(Ebin)}$, as described in sec. 3.4. Table 3 provides the IR counterparts from the AllWISE catalogue (Cutri et al., 2021) which are meant to help guide the optical observations [^12] For 4FGLJ0351.2-6103, the AllWISE IR counterpart is tentative; and for 4FGLJ2127.6-5959, a tentative IR counterpart is close by NGC7059..
[…]
Name R.A.(deg) Dec.(deg) 4FGLdr4 ASSOC1-4FGL TS $\rm E^{bin}_{max}$ Rmag AllWISE
4FGLJ1634.9+1222 248.76049 12.3619 J1634.9+1222 NVSSJ163502+122142 15 233 20.04 J163502.41+122142.1
3HSPJ013632.6+390559 24.13579 39.0997 J0136.5+3906 B30133+388 4708 183 15.86 J013632.59+390559.1
3HSPJ130738.0--425938 196.90825 --42.9941 J1307.6--4259 1RXSJ130737.8--425940 732 150 16.26 J130737.98--425938.9
4FGLJ0840.1--0225 130.06121 --2.4531 J0840.1--0225 PMNJ0840--0227 37 149 14.57 J084014.69--022711.3
4FGLJ1334.1--3521 203.55010 --35.3372 J1334.1--3521 PKS1331--350 21 138 12.25 J133412.03--352014.2
3HSPJ074642.0--475455 116.67629 --47.9152 J0746.6--4754 PMNJ0746--4755 1223 134 18.18 J074642.31--475455.0
3HSPJ230436.7+370507 346.15295 37.0854 J2304.6+3704 1RXSJ230437.1+370506 297 98 20.2 J230436.71+370507.4
3HSPJ102634.4--854314 156.64316 --85.7206 J1027.0--8542 PKS1029--85 528 95 17.27 J102634.36--854314.2
5BZBJ0700--6610 105.13017 --66.1792 J0700.5--6610 PKS0700--661 986 94 15.29 J070031.25--661045.2
3HSPJ230940.8--363248 347.42012 --36.5468 J2309.7--3632 WISEJ230940.84--363248 122 87 18.62 J230940.84--363248.7
3HSPJ033349.0+291631 53.45416 29.2754 J0333.7+2916 TXS0330+291 591 85 19.1 J033349.00+291631.5
5BZBJ1849--4314 282.358 --43.2369 J1849.4--4313 PMNJ1849--4314 121 82 17.8 J184925.92--431413.3
4FGLJ0703.2--3914 105.80268 --39.2385 J0703.2--3914 1RXSJ070312.7--391417 52 81 17.6 J070312.65--391418.8
4FGLJ1608.0--2038 241.98719 --20.6617 J1608.0--2038 NVSSJ160756--203942 50 80 18.22 J160756.90--203943.5
3HSPJ130421.0--435310 196.0875 --43.8861 J1304.3--4353 1RXSJ130421.2--435308 1049 77 17.48 J130421.00--435310.2
4FGLJ2345.9+3413 356.47985 34.2342 J2345.9+3413 1RXSJ234554.2+341419 21 74 17.8 J234555.21+341404.5
4FGLJ2143.9+3337 325.95889 33.6196 J2143.9+3337 MG3J214351+3337 74 73 20.6 *J214350.20+333711.7
4FGLJ1345.6--3356 206.42935 --33.9453 J1345.6--3356 NVSSJ134543--335643 76 73 19.58 J134543.04--335643.3
5BZBJ2108--6637 317.21592 --66.6229 J2108.9--6638 PKS2104--668 198 73 18.22 J210851.80--663722.7
4FGLJ0836.0--8015 128.95466 --80.2696 J0836.0--8015 2MASSJ08354940--8016114 29 69 20.74 J083549.36--801610.8
4FGLJ2300.8--0736 345.22798 --7.5954 J2300.8--0736 2MASSJ23005469--0735438 34 69 19.51 J230054.70--073543.5
3HSPJ081012.0--703047 122.55016 --70.5130 J0809.9--7028 SUMSSJ081011--703048 12 68 19.44 J081012.04--703047.1
3HSPJ104756.9--373730 161.98725 --37.6252 J1047.9--3738 GALEXASCJ104756--373730 45 65 19.7 J104756.94--373730.8
4FGLJ1804.4+5249 271.09484 52.8205 J1804.4+5249 6CB180317.2+524912 14 65 15.87 J180422.64+524915.2
3HSPJ195547.9+021512 298.94941 2.2535 J1955.7+0214 NVSSJ195547+021514 92 63 20.95 J195547.95+021517.9
4FGLJ1836.5+1948 279.13380 19.8461 J1836.5+1948 NVSSJ183632+195047 54 62 18.83 J183632.11+195046.3
5BZBJ0718--4319 109.68183 --43.3304 J0718.6--4319 PMNJ0718--4319 760 62 17.73 J071843.63--431949.7
4FGLJ2025.3--2231 306.31322 --22.5051 J2025.3--2231 NVSSJ202515--223016 55 62 18.15 J202515.17--223018.4
4FGLJ2127.6--5959 321.87039 --60.0137 J2127.6--5959 NGC7059(?) 16 60 17.79 *J212729.34--600055.8
4FGLJ0624.2--2943 96.09291 --29.7469 J0624.2--2943 1RXSJ062422.3--294449 19 59 21.46 J062422.49--294446.7
4FGLJ2049.0--4020 312.29171 --40.3419 J2049.0--4020 PKS2045--405 19 59 19.53 J204910.02--402030.9
3HSPJ151444.0--772254 228.68341 --77.3817 J1514.4--7719 1RXSJ151448.8--772249 30 58 19.74 J151444.02--772254.2
4FGLJ0351.2--6103 57.75249 --61.0467 J0351.2--6103 SUMSSJ035100--610248 15 57 16.61 *J035059.83--610241
4FGLJ0928.7--3529 142.20762 --35.4969 J0928.7--3529 NVSSJ092849--352947 15 56 19.64 J092849.82--352948.8
3HSPJ145543.7--760052 223.93204 --76.0145 J1455.8--7601 SUMSSJ145543--760054 45 55 18.41 J145543.69--760052.2
4FGLJ0942.7--2028 145.65761 --20.4536 J0942.7--2028 1RXSJ094237.9--202720 75 55 19.96 J094237.81--202713.2
3HSPJ094709.5--254100 146.78966 --25.6833 J0947.1--2541 1RXSJ094709.2--254056 397 55 18.16 J094709.52--254059.9
4FGLJ1513.4--3721 228.32775 --37.3365 J1513.4--3721 2MASSJ15131867--3720114 56 54 17.83 J151318.66--372011.5
3HSPJ080215.9--094210 120.56625 --9.7030 J0802.3--0942 WISEJ080215.63--0942(?) 235 54 19.04 J080215.90--094210.9
4FGLJ1828.0+2634 276.98555 26.5535 J1828.0+2634 NVSSJ182756+263313 40 50 19.4 J182756.54+263313.2
3HSPJ213004.8--563222 322.51987 --56.5394 -- -- 18 50 20.48 J182756.54+263313.2
3HSPJ134706.9--295842 206.77866 --29.9784 J1347.1--2959 NVSSJ134706--295840 62 50 18.8 J134706.88--295842.4
Table 4: This table lists the sources in the catalogue with uncertain redshift information, ordered by their $\tau_{(E,z)}$ values calculated at $\rm E^{bin}_{max}$. The first columns show the source names, R.A., and Dec. in degrees (J2000), respectively. The fourth column presents redshift information from the literature, with a ‘(?)’ flag indicating uncertain or photometric values and a ‘>’ symbol denoting lower limits. The ‘z-origin’ column specifies the literature reference for the redshift, using the short names described in Section 3.1. The ‘b(deg)’ column shows the galactic latitude in degrees, followed by the associated ‘4FGL-DR4’ name. The ‘$\rm TS^{1CGH}_{10GeV}$’ column provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations. The columns ‘$\rm E_{max}^{bin}$’ and ‘$\rm\tau_{Ebin}$’ list the largest energy bin detected and the corresponding EBL optical depth $\tau_{(E,z)}$, respectively. For sources whose Name’ begins with 4FGLJ, the R.A. and Dec. correspond to the associated counterpart position(see the ‘ASSOC1-4FGL’ column in, Ballet et al., 2023).
[…]
3HSPJ224753.2+441315 341.97175 44.22097 1.9(?) Foschini22 --13.2 J2247.8+4413 375 179 5.08
3HSPJ015307.4+751742 28.28070 75.29522 2.35(?) Foschini22 12.9 J0153.0+7517 168 149 4.94
4FGLJ0400.7+3920 60.18911 39.35271 1.1(?) Foschini22 --10.2 J0400.7+3920 18 209 3.34
3HSPJ200506.0+700439 301.27487 70.07763 2.32(?) Foschini22 19.4 J2005.1+7003 311 66 1.60
3HSPJ052542.4--601340 81.42675 --60.22783 1.78(?) Kaur16 --33.8 J0525.6--6013 249 70 1.21
4FGLJ1937.0+8354 294.41522 83.94139 1.94(?) Foschini22 25.7 J1937.0+8354 39 61 1.12
4FGLJ1858.1+7318 284.58486 73.28700 0.471(?) Foschini22 25.4 J1858.1+7318 43 221 1.10
3HSPJ031612.7+090443 49.05304 9.07866 0.372(?) 3HSPcat --39.5 J0316.2+0905 621 255 0.99
3HSPJ142829.9+743002 217.12454 74.50061 0.31(?) 3HSPcat 41.0 J1428.8+7429 18 294 0.95
3HSPJ144800.6+360831 222.00245 36.142 >0.738 Paiano20 63.7 J1448.0+3608 1085 137 0.94
5BZBJ1248+5820 192.07825 58.34131 0.508(?) Foschini22 58.7 J1248.3+5820 4149 189 0.94
5BZBJ0007+4712 1.99988 47.20214 >1.659 Shaw13 --15.0 J0008.0+4711 709 61 0.87
3HSPJ122337.0--303250 185.90420 --30.54725 >0.875 Desai19 31.9 J1223.6--3032 76 113 0.87
4FGLJ0318.7+2135 49.69028 21.57685 1.83(?) Foschini22 --29.6 J0318.7+2135 171 55 0.85
3HSPJ063059.5--240646 97.748 --24.11280 >1.239 3HSPcat --14.9 J0630.9--2406 3589 78 0.80
5BZBJ1918--4111 289.56687 --41.19189 >1.591 Shaw13 --22.2 J1918.2--4111 555 60 0.79
5BZBJ1925--2219 291.41579 --22.32642 1.35(?) Foschini22 --17.1 J1925.8--2220 37 69 0.76
5BZBJ2300+3137 345.09512 31.61789 >1.498 Shaw13 --25.5 J2300.3+3136 283 62 0.76
5BZBJ1314+2348 198.68254 23.80744 2.053(?) SDSSdr18 83.7 J1314.7+2348 387 46 0.76
4FGLJ0453.1+6322 73.30189 63.35496 2.1(?) Foschini22 12.1 J0453.1+6322 39 45 0.75
3HSPJ154015.1+815505 235.06625 81.91822 >0.67 Paiano17 32.9 J1540.1+8155 938 129 0.72
3C 66A 35.67334 43.04319 >0.3347 Furniss13 --16.7 J0222.6+4302 7060 230 0.70
5BZBJ0856--1105 134.17421 --11.08733 >1.397 Shaw13 21.4 J0856.6--1105 517 64 0.70
5BZBJ0612+4122 93.21325 41.37706 >1.107 Paiano20 10.9 J0612.8+4122 1383 80 0.68
3HSPJ181118.0+034113 272.82508 3.68711 0.717(?) Foschini22 10.6 J1811.3+0340 473 117 0.67
3HSPJ043145.1+740326 67.93775 74.05738 1.35(?) Foschini22 17.3 J0431.8+7403 226 65 0.67
3HSPJ103744.3+571155 159.43458 57.19880 >0.62 Meisner10 51.7 J1037.7+5711 3071 131 0.65
3HSPJ003514.1+151504 8.81129 15.25116 >0.64 3HSPcat --47.4 J0035.2+1514 485 122 0.60
3HSPJ112048.1+421212 170.20025 42.2035 >0.35 3HSPcat 66.1 J1120.8+4212 1811 199 0.58
S5 0716+714 110.47250 71.34332 0.31(?) Foschini22 28.0 J0721.9+7120 11282 218 0.57
3HSPJ170433.8--052840 256.14095 --5.47797 >0.7 Paiano17 20.7 J1704.5--0527 244 109 0.55
3HSPJ152048.9--034851 230.20370 --3.81433 >0.868 Goldini20 42.5 J1520.8--0348 421 88 0.53
3HSPJ193419.6+600139 293.58175 60.02763 1.38(?) Foschini22 18.1 J1934.2+6002 44 56 0.52
4FGLJ1304.2--2412 196.06955 --24.20465 1.26(?) Foschini22 38.5 J1304.2--2412 155 60 0.51
5BZBJ1312--2156 198.1315 --21.93983 >1.485 Shaw13 40.6 J1312.4--2156 358 50 0.50
4FGLJ0028.5+2001 7.12424 20.00742 1.552(?) 4LACdr3 --42.5 J0028.5+2001 489 48 0.50
3HSPJ175615.1+552218 269.06625 55.37166 >0.657 3HSPcat 29.7 J1756.3+5522 264 106 0.47
4FGLJ0355.3+3909 58.81912 39.15272 0.846(?) Foschini22 --11.0 J0355.3+3909 18 85 0.47
3HSPJ210421.9--021238 316.09137 --2.21080 >0.45 3HSPcat --30.3 J2104.3--0212 88 140 0.43
3HSPJ222129.3--522527 335.37208 --52.4243 >0.34 3HSPcat --52.3 J2221.5--5225 426 174 0.42
3HSPJ125359.3+624257 193.49725 62.716 >0.867 Shaw13 54.4 J1253.8+6242 179 79 0.42
4FGLJ1537.9--1344 234.48793 --13.7262 0.984(?) Foschini22 32.5 J1537.9--1344 31 70 0.42
5BZBJ2016--0903 304.10013 --9.0592 >0.605 Shaw13 --22.9 J2016.3--0903 286 108 0.42

### v3

@@
 keywords:
-gamma-rays: galaxies, catalogues – radiation mechanisms: non-thermal
+gamma-rays: galaxies – galaxies: active – catalogues – radiation mechanisms: non-thermal
 pubyear: 20252024pagerange: Mapping the Cosmic Gamma-ray Horizon: The 1CGH Catalogue of Fermi-LAT detections above 10 GeVB
@@
 The attenuation of $\gamma$-rays by the EBL defines a statistical boundary for the farthest distance at which VHE photons of a given energy are likely to be detected, known as the Cosmic Gamma-ray Horizon (CGH).
-The concept of the CGH was initially explored by Nikishov (1961); Gould & Schréder (1967), while Fazio & Stecker (1970) -with the Fazio Stecker relation- advanced the discussion of $\gamma$-ray attenuation in context of a cosmological origin for the isotropic $\gamma$-ray background.
+The CGH concept was initially explored by Nikishov (1961); Gould & Schréder (1967), while Fazio & Stecker (1970) -with the Fazio Stecker relation- advanced the discussion of $\gamma$-ray attenuation in context of a cosmological origin for the isotropic $\gamma$-ray background.
 Interestingly, the VHE spectra of distant blazars can exhibit an unexpected hardening at the highest energies (e.g.
@@
 To create a sample of sky positions and define our $\gamma$-ray targets, we combined several major blazar and AGN catalogues.
-We began by merging the ‘5th edition of the Roma-BZCAT Multi-Frequency Blazar Catalog’ (5BZcat, Massaro et al., 2015a) with the ‘3rd edition of the High Synchrotron Peak sample’ (3HSP, Chang et al., 2019).
+We began by merging the ‘5th edition of the Roma-BZCAT Multi-Frequency Blazar Catalog’ (5BZcat; Massaro et al., 2015a) with the ‘3rd edition of the High Synchrotron Peak sample’ (3HSP; Chang et al., 2019).
 These catalogues were chosen for their extensive blazar coverage and reliability.
-Both 5BZcat and 3HSP are advanced versions of multifrequency catalogues containing thousands of blazars and blazar candidates.
-These catalogues have been iteratively refined over time (e.g., refer to BZcat, 1WHSP, and 2WHSP catalogues for earlier versions Massaro et al., 2009; Arsioli et al., 2015; Chang et al., 2017), and represent significant repositories for the very high-energy community.
+Both 5BZcat and 3HSP are advanced versions of multifrequency catalogues containing thousands of blazars and blazar candidates, and represent significant repositories for the VHE community.
+These catalogues have been iteratively refined over time, building upon the initial BZcat (Massaro et al., 2009), and the First and Second Catalogues of High Synchrotron Peaked Blazars and Candidates, 1WHSP and 2WHSP (Arsioli et al., 2015; Chang et al., 2017).
 We also included the latest version (V3.4) of TeVcat (Wakely & Horan, 2008), which contains a follow-up list of VHE-detected sources; http://tevcat.uchicago.edu.
@@
 (2023, 2024).
-We further include all sources classified as AGNs, blazars, and blazar candidates in the ‘Fourth Catalog of Active Galactic Nuclei Detected by the Fermi Large Area Telescope: Data Release 3’ (4LAC-DR3, Ajello et al., 2022), since many of those candidates are not covered by the 5BZcat and 3HSP catalogues.
+We further include all sources classified as AGNs, blazars, and blazar candidates in the ‘Fourth Catalog of Active Galactic Nuclei Detected by the Fermi Large Area Telescope: Data Release 3’ (4LAC-DR3; Ajello et al., 2022), since many of those candidates are not covered by the 5BZcat and 3HSP catalogues.
 To remove duplicate entries, we use TopCat (Taylor, 2005) for internal matching; For the 5BZcat and 3HSP samples, the R.A.
@@
 For TeVcat and 4LAC-DR3 sources, we used the associated counterpart position whenever available (e.g., radio, infrared, or optical), as these typically offer better astrometric precision compared to the $\gamma$-ray position.
-With this sample of unique sources, we searched for $\gamma$-ray associations in the 4FGL-DR4 catalogue (Ballet et al., 2023) and the 3FHL catalogue (Ajello et al., 2017), considering the 95% containment region reported for each source.
+With this sample of unique sources, we searched for $\gamma$-ray associations in the ‘Fermi LAT 14-Year Point Source Catalogue’ (4FGL-DR4; Ballet et al., 2023) and the 3FHL catalogue (Ajello et al., 2017), considering the 95% containment region reported for each source.
 At this point, we have successfully listed unique seed positions recovered from the 5BZcat, 3HSP, TeVcat and 4LAC-DR3 catalogues; But still, many 4FGL-DR4 high-latitude sources ($|b|>10^{\circ}$) were not included in our initial sample of 10 GeV candidates.
@@
 In our analysis, we adopt a low-energy threshold of 10 GeV –similar to the 3FHL catalogue– due to the improved angular resolution and lower background contamination compared to lower thresholds (Ajello et al., 2017).
-For the upper energy limit, we set a conservative threshold of 800 GeV, whereas 3FHL extended the analysis up to 2 TeV.
-This decision follows from the discussion in the 4FGL paper (Section 3.2 Abdollahi et al., 2020), which highlights that a broadband analysis reaching 1 TeV can introduce uncertainties in the energy flux estimate, particularly for hard-spectrum sources.
+For the upper energy limit, we set a conservative threshold of 800 GeV, whereas 3FHL extended the analysis to 2 TeV.
+This decision follows from the discussion in the 4FGL paper (Section 3.2 of Abdollahi et al., 2020), which highlights that a broadband analysis reaching 1 TeV can introduce uncertainties in the energy flux estimate, particularly for hard-spectrum sources.
 Therefore, our choice for the high-energy threshold is meant to mitigate uncertainties regarding the spectral fitting of all sources in the ROI region.
 2.3 Detection Threshold for Multifrequency-Selected Targets
-The use of multifrequency seed positions to detect $\gamma$-ray sources has been successfully applied in numerous works (Arsioli & Chang, 2017; Arsioli et al., 2018; Arsioli & Polenta, 2018a; Arsioli et al., 2020) and has proven effective in uncovering extreme blazars close to the Fermi-LAT detection threshold.
+The use of multifrequency seed positions to detect $\gamma$-ray sources has been successfully applied in numerous works (Arsioli & Chang, 2017; Arsioli et al., 2018; Arsioli & Polenta, 2018; Arsioli et al., 2020) and has proven effective in uncovering extreme blazars close to the Fermi-LAT detection threshold.
 Using this approach, the analysis for each sky position is independent, investigating a single $\gamma$-ray candidate at a time.
@@
 […]
-In addition, this relatively low selection threshold aims not only to highlight faint sources but also to capture those that undergo short-lived flare episodes, which might otherwise be missed over a long integration period due to signal dilution (Arsioli & Polenta, 2018b).
+In addition, this relatively low selection threshold aims not only to highlight faint sources but also to capture those that undergo short-lived flare episodes, which might otherwise be missed over a long integration period due to signal dilution (Arsioli & Polenta, 2018).
 As the true signal (i.e.
@@
 Given that 8200 seed positions were analysed, the expected number of spurious detections in the pre-selection phase is approximately $\rm 8200\times 0.0025=20.5$.
-This value represents an upper limit on contamination.
+This value represents an upper limit of contamination.
 In next section, we discuss an additional selection criteria –beyond the selection based on TS level alone– to avoid spurious detections and improve reliability of the final catalogue.
@@
 The PSF 68% containment radius varies by energy and event type.
-For instance, at 30 GeV, the containment radius for PSF1 events is approximately 0.12, as described in the PASS8 documentation [^4] Fermi-LAT PASS8 PSF: https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm.
+For example, at 30 GeV, the containment radius for PSF1 events is approximately 0.12, as described in the Pass 8 documentation [^4] Fermi-LAT Pass 8 PSF: https://www.slac.stanford.edu/exp/glast/groups/canda/lat_Performance.htm.
 Using the gtselect routine, we created a subsample of the all-sky Fermi-LAT data, selecting only events with energy greater than 10 GeV.
@@
 Table 1 presents a selection of newly detected sources with TS above 16, including the seed-source name, right ascension (R.A.), declination (Dec.), redshift, and power-law model parameters associated with each new $\gamma$-ray detection [^7] The full version of the First Cosmic Gamma-ray Horizon catalogue will be available in public data repositories such as Vizier https://vizier.cds.unistra.fr/viz-bin/VizieR and Author’s GitHub..
-Figure 1: The distribution of detection significance ($\sigma$) for the 1CGH and 3FHL samples (respectively, blue and red-dashed lines).
+Figure 1: The distribution of detection significance ($\sigma$) for the 1CGH and 3FHL samples (respectively, blue line and red-dashed line).
 In Fig.
-1 we show the distribution of detection significance for the entire 1CHG and 3FHL catalogues.
+1 we show the distribution of the detection significance for the entire 1CHG and 3FHL catalogues.
 This representation is meant as a qualitative overview on both samples, keeping in mind that the 1CGH is focus on seed-positions of blazars and blazar-candidates, while the 3FHL includes detections of all source types.
@@
 following a trend similar to 3FHL), and includes an extra bin covering the 3<$\sigma$<4 range.
-We should note that, in the full 1CGH catalogue, 26 sources [^8] Comprising 3 unassociated sources, 4 BL Lacs, 8 Flat Spectrum Radio Quasars (FSRQs), and 11 blazars of unknown type.
+We should note that in the full 1CGH catalogue, 26 sources [^8] Comprising 3 unassociated sources, 4 BL Lacertae, 8 Flat Spectrum Radio Quasars (FSRQs), and 11 blazars of unknown type.
 reached the spectral index limit of $\Gamma=6.5$ constrained by our likelihood analysis setup (see Section 2.2).
@@
 For these 53 (26+27) cases, the derived fluxes should be interpreted as upper limits.
-While this affects a small fraction of the sample, the vast majority of sources exhibit well-constrained spectral fits, reinforcing the robustness of the 1CGH catalogue for high-energy studies.
+While this affects only a small fraction of the sample, the vast majority of sources exhibit well-constrained spectral fits, reinforcing the robustness of the 1CGH catalogue for high-energy studies.
+Figure 2: The photon index ($\Gamma$) versus the integral flux in the 10 GeV to 800 GeV energy band.
+The green points represent 1CGH sources with counterparts in both 3FHL and 4FGL-DR4 catalogues; the blue open diamonds represents 1CGH sources with counterparts in 4FGL-DR4 only; and the magenta circles represents new $\gamma$-ray identifications.
+In Fig.
+2 we present the photon index ($\Gamma$) versus the integrated flux derived from a power-law fit over the 10–800 GeV energy range.
+The plot shows that the 1CGH sample extends the 3FHL coverage toward fainter fluxes, effectively lowering the sensitivity limit for detections above 10 GeV.
+As in the 3FHL catalogue, the detection threshold in the 1CGH sample shows little dependence on photon index, with the new identifications clustering at the lowest fluxes.
+Moreover, the overlap between the 1CGH and 4FGL sources highlights a robust improvement in completeness above 10 GeV, which is complemented by the newly identified sources.
 3.1 Extensive Redshift Review
@@
 The numerous dedicated optical observation campaigns highlight the ongoing high demand for the characterization of blazars and blazar candidates, particularly in connection with their $\gamma$-ray counterparts.
+Starting from the redshift information available in the reference catalogues (i.e., 5BZcat, 3HSP, 4LAC-DR3, and TeVcat), our literature review allowed us to assign or update redshift values for 967 1CGH sources –including 377, 437, and 153 sources with z-flags 1, 2, and 3, respectively.
+Of the total 2791 1CGH sources, 1062 (38.2%) have spectroscopically measured redshift values, 855 (30.6%) have uncertain/photometric values, 210 (7.5%) have lower-limit values, and 664 (23.7%) have no available redshift information.
 In Figure 3, we show the redshift distribution for the 1CGH subsample with z-flag=1 (i.e.
@@
 It highlights the significance of reaching an extensive redshift description of the entire 1CGH sample and aims to further motivate the community’s ongoing efforts in the optical characterization of these rare sources (see Section 3.1).
-Here we should draw attention to the number of 1CGH sources which lack a proper redshift characterization: Of the total 2791 1CGH sources, 1062 (38.2%) have spectroscopically measured redshift values, 855 (30.6%) have uncertain/photometric values, 210 (7.5%) have lower limit values, and 664 (23.7%) have no available redshift information.
-Therefore, 1729 ($\sim$72%) of the 1CGH catalogue would benefit from dedicated observation campaigns to improve the overall redshift description.
+Here we draw attention to the number of 1CGH sources that lack robust redshift characterization; As mentioned in Section 3.1, a total of 1729 1CGH sources ($\sim$72%) are assigned redshift values that are either uncertain, lower limits, or are absent.
+All of those cases would benefit from dedicated observational campaigns to improve the overall redshift description, therefore representing a significant challenge for CGH studies and the very high-energy community as a whole.
 3.3 Identification of Sources with Detectable EBL Absorption
@@
 (2021) as a reference framework to identify sources where gamma-ray absorption might be detectable.
-The Saldana-Lopez model offers an empirical determination of the evolving EBL spectrum and its uncertainties up to $\sim$6; it is based on galaxy counts and multifrequency data (from UV to far-IR) covering the Hubble Space Telescope’s CANDELS survey (Grogin et al., 2011), and was designed to minimise uncertainties especially at higher redshift.
-Our ultimate goal, however, is to highlight a sample in which $\tau_{(E,z)}$ values can be derived from the data, independently of any specific EBL model.
+The Saldana-Lopez model offers an empirical determination of the evolving EBL spectrum and its uncertainties up to $\sim$6; it is based on galaxy counts and multifrequency data (from UV to far-IR) covering the Hubble Space Telescope’s ‘Cosmic Assembly Near-infrared Deep Extragalactic Legacy Survey’ (CANDELS), and was designed to minimise uncertainties especially at higher redshift (Grogin et al., 2011; Koekemoer et al., 2011).
+However, our ultimate goal is to highlight a sample in which the $\tau_{(E,z)}$ values can be derived from the data, independently of any specific EBL model.
 The estimates of Saldana-Lopez et al.
@@
 (2010) and both are representative of the most severe attenuation models among the ones considered.
-We use an EBL optical depth of $\rm\tau_{(E,z)}=0.1$ as a lower limit to introduce an absorption flag (ABS-flag), indicating whether spectral absorption might be detectable at the highest energy bins.
+We used an EBL optical depth of $\rm\tau_{(E,z)}=0.1$ as the lower limit to introduce an absorption flag (ABS-flag), indicating whether spectral absorption might be detectable at the highest energy bins.
 For cases where $\rm E^{bin}_{max}$ lies in the $\rm\tau_{(E,z)}>0.1$ regime, the largest energy bin observed with Fermi-LAT likely experiences a detectable absorption greater than 10% (i.e., >($\rm 1-e^{-0.1}$)).
-These sources can be used to investigate the opacity along redshift, and have an ABS-flag set to ‘1’.
+These sources can be used to investigate the opacity along redshift, and have ABS-flag set to ‘1’.
 For sources with $\rm E^{bin}_{max}$ in the $\rm\tau_{(E,z)}<0.1$ regime, the Fermi-LAT spectrum is likely unaffected by EBL attenuation, and these sources have an ABS-flag set to ‘0’ in the 1CGH catalogue.
@@
 The $\rm\tau_{(E,z)}$ values –calculated using the Saldana-Lopez EBL model at $\rm E^{bin}_{max}$ Saldana-Lopez et al.
-(2021)– are listed under the column ‘$\rm\tau_{Ebin}$’.
-Additionally, the absorption fraction ($\rm 1-e^{-\tau_{(E,z)}}$) is recorded in the ‘$\rm Abs_{Ebin}$’ column, which can be used to create subsamples based on different absorption thresholds.
-Notably, for sources with multiple energy bins experiencing detectable absorption, a single source can provide $\rm\tau_{(E,z)}$ estimates for multiple energy levels at a given redshift.
+(2021)– are listed in the column ‘$\rm\tau_{Ebin}$’.
+Furthermore, the absorption fraction ($\rm 1-e^{-\tau_{(E,z)}}$) is recorded in the ‘$\rm Abs_{Ebin}$’ column, which can be used to create subsamples based on different absorption thresholds.
+In particular, for sources with multiple energy bins that experience detectable absorption, a single source can provide $\rm\tau_{(E,z)}$ estimates for multiple energy levels at a given redshift.
 3.4 Best Candidates for Optical Observation Campaigns
@@
 In this work, we presented the First Cosmic Gamma-ray Horizon (1CGH) catalogue, which lists all-sky Fermi-LAT detections above 10 GeV, compiled from 16 years of observations.
-The catalogue includes 2791 sources, with 62 representing new $\gamma$-ray detections, and to date, it represents an advanced sample to study the transparency of the universe to VHE photons.
+The catalogue includes 2791 sources, with 62 representing new $\gamma$-ray detections, and to date, it represents a valuable sample to study the transparency of the Universe to VHE photons.
 We significantly improved the redshift characterization of our catalogue by incorporating new spectroscopic redshift estimates whenever possible.
-By meticulously reviewing nearly 70 dedicated observational campaigns, we reduced a substantial portion of the redshift knowledge-gap, categorising redshift reliability through our z-flag system.
-The 1CGH catalogue highlights sources experiencing varying degrees of $\gamma$-ray attenuation, and will contribute to a better determination of the EBL’s density and its evolution over cosmic time.
+By meticulously reviewing nearly 70 dedicated observational campaigns, we significantly reduced the redshift knowledge gap, introducing a z-flag system to categorise redshift reliability.
+The 1CGH catalogue highlights sources experiencing varying degrees of $\gamma$-ray attenuation and will contribute to a better determination of the EBL’s density and its evolution over cosmic time.
 Using the Saldana-Lopez EBL model as a reference framework, we identified sources with $\gamma$-ray absorption above the $\rm\tau_{(E,z)}>0.1$ limit, creating a targeted sample where absorption is likely measurable.
-In addition, we introduced a complementary data-exploration method by calculating the mean energy of the four highest-energy photons $\rm E^{bin}_{max}$, rather than focusing solely on the highest-energy event.
+In addition, we introduced a complementary data exploration method by calculating the mean energy of the four highest-energy photons $\rm E^{bin}_{max}$, rather than relying solely on the highest-energy event.
 This provides a more stable and statistically reliable measure of the characteristic ‘largest energy level’ observed for each source.
-We have identified sources that should be prioritised for follow-up optical observation campaigns.
-These include: (i) sources that lack redshift information but have a clear optical or radio association, prioritised based on the highest $\rm E^{bin}_{max}$ values; and (ii) sources with lower limit or uncertain redshift estimates that could benefit from further spectroscopic exploration, prioritised by their current $\tau$ values at $\rm E^{bin}_{max}$.
+We identified sources that should be prioritised for follow-up optical observation campaigns.
+These include: (i) sources that lack redshift information but have a clear optical or radio association, prioritised based on the highest $\rm E^{bin}_{max}$ values; and (ii) sources with lower limit or uncertain redshift estimates that could benefit from further spectroscopic exploration, prioritised according to their $\tau$ values measured at $\rm E^{bin}_{max}$.
 These targets will not only help refine the redshift description of the catalogue but also improve the accuracy of EBL measurements at higher redshift, thus reducing uncertainties in SFR density determination during the EoR, at $\sim$6-7.
+Following our redshift review (Section 3.1), approximately 72% of the 1CGH sources still lack robust redshift determination.
+This significant gap poses challenges for the very high-energy community as a whole, and calls attention to the need for extensive follow-up optical observation campaigns targeting $\gamma$-ray blazars.
 Acknowledgements
@@
 Appendix A Column Descriptions of the 1CGH Catalogue
-The 1CGH catalogue contains detailed metadata for 2791 sources detected above 10 GeV based on 16 years of Fermi-LAT observations. Table 2 provides a description of each column, detailing its content.
+The 1CGH catalogue contains detailed metadata for 2791 sources detected above 10 GeV based on 16 years of Fermi-LAT observations. Table 2 provides a description of each column and details its content.
 Table 2: Column descriptions of the 1CGH catalogue. The table lists the column names, units where applicable, and detailed explanations of their contents.
@@
 z_3HSP Redshift from the 3HSP catalogue, if available.
-zflag_3HSP Redshift flag from the 3HSP catalogue.
+zflag_3HSP Redshift flag from the 3HSP catalogue: (1) spectroscopic, (2) uncertain, (3) lower limit redshift, (4) photometric -with featureless optical spectrum, (5) photometric -without optical spectrum.
 4FGL_counter_name Counterpart name in 4FGL-DR4.
@@
 3.4.
-Table 3 provides the IR counterparts from the AllWISE catalogue (Cutri et al., 2021) which are meant to help guide the optical observations [^12] For 4FGLJ0351.2-6103, the AllWISE IR counterpart is tentative; and for 4FGLJ2127.6-5959, a tentative IR counterpart is close by NGC7059..
-[…]
-Name R.A.(deg) Dec.(deg) 4FGLdr4 ASSOC1-4FGL TS $\rm E^{bin}_{max}$ Rmag AllWISE
+Table 3 provides the IR counterparts from the ‘Wide-field Infrared Survey Explorer’ AllWISE catalogue (Cutri et al., 2013, 2021) which are meant to help guide optical observations [^12] For 4FGLJ0351.2-6103, the AllWISE IR counterpart is tentative; and for 4FGLJ2127.6-5959, a tentative IR counterpart is close by NGC7059..
+Table 3: This table lists the sources in the catalogue without redshift information, ordered by their $\rm E^{bin}_{max}$, and where we could retrieve clear IR/Optical counterparts for observation.
+The first three columns show the source names, right ascension (R.A.), and declination (Dec.) in degrees (J2000), respectively.
+The column ‘4FGL-DR4’ and ‘ASSOC1-4FGL’ shows the source and counterpart names reported in 4FGL-DR4 catalogue.
+Next, the column ‘TS’ provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations.
+The column ‘$\rm E_{max}^{bin}$’ list the highest energy bin detected.
+The column ‘Rmag’ provides the R band magnitude retrieved using SSDC Sky Explorer.(https://tools.ssdc.asi.it/).
+The last column, ‘AllWISE’ provides the IR counterpart from the AllWISE catalogue (cases marked with a ‘*’ sign represent a tentative association and should be considered with care).
+For sources whose ’Name’ begins with 4FGLJ, the R.A.
+and Dec.
+correspond to the counterpart position (‘ASSOC1’, Ballet et al., 2023).
+Name R.A.(deg) Dec.(deg) 4FGL-DR4 ASSOC1-4FGL TS $\rm E^{bin}_{max}$ Rmag AllWISE
 4FGLJ1634.9+1222 248.76049 12.3619 J1634.9+1222 NVSSJ163502+122142 15 233 20.04 J163502.41+122142.1
@@
 The ‘b(deg)’ column shows the galactic latitude in degrees, followed by the associated ‘4FGL-DR4’ name.
-The ‘$\rm TS^{1CGH}_{10GeV}$’ column provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations.
+The ‘TS’ column provides the Test Statistic value derived for 1CGH sources detected above 10 GeV, integrating over 16 years of Fermi-LAT observations.
 The columns ‘$\rm E_{max}^{bin}$’ and ‘$\rm\tau_{Ebin}$’ list the largest energy bin detected and the corresponding EBL optical depth $\tau_{(E,z)}$, respectively.
@@
 correspond to the associated counterpart position(see the ‘ASSOC1-4FGL’ column in, Ballet et al., 2023).
-[…]
+Name R.A. (deg) Dec. (deg) z z-origin b(deg) 4FGL-DR4 TS $\rm E^{bin}_{max}$ $\rm\tau_{Ebin}$
 3HSPJ224753.2+441315 341.97175 44.22097 1.9(?) Foschini22 --13.2 J2247.8+4413 375 179 5.08